# OpeNLGauge: An Explainable Metric for NLG Evaluation with Open-Weights LLMs

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated great potential as evaluators of NLG systems, allowing for high-quality, reference-free, and multi-aspect assessments. However, existing LLM-based metrics suffer from two major drawbacks: reliance on proprietary models to generate training data or perform evaluations, and a lack of fine-grained, explanatory feedback. In this paper, we introduce OpeNLGauge, a fully open-source, reference-free NLG evaluation metric that provides accurate explanations based on error spans. OpeNLGauge is available as a two-stage ensemble of larger open-weight LLMs, or as a small fine-tuned evaluation model, with confirmed generalizability to unseen tasks, domains and aspects. Our extensive meta-evaluation shows that OpeNLGauge achieves competitive correlation with human judgments, outperforming state-of-the-art models on certain tasks while maintaining full reproducibility and providing explanations more than twice as accurate.
@@
Ivan Kartáč  and  Mateusz Lango  and  Ondřej Dušek Charles University, Faculty of Mathematics and Physics Institute of Formal and Applied Linguistics Prague, Czech Republic ivan.kartac@gmail.com, {lango,odusek}@ufal.mff.cuni.cz
@@ 1 Introduction
Evaluating Natural Language Generation (NLG) systems remains a challenging research problem. Traditional overlap-based metrics, such as BLEU Papineni et al. (2002a) and ROUGE Lin (2004a), are still widely used but exhibit limited correlation with human judgments, particularly when assessing modern NLG systems Novikova et al. (2017a). With the rise of pre-trained language models, the research community began to shift toward model-based metrics that better capture semantic similarity, yet their performance was still unsatisfactory Yan et al. (2023); Glushkova et al. (2023).
Figure 1: Overview of the proposed approach: OpeNLGaugeens ensemble and its distilled version OpeNLGaugeft.
@@
Recently, Large Language Models (LLMs) have demonstrated remarkable potential in imitating human evaluation of generated text Jiang et al. (2023); Xu et al. (2023a); Hu et al. (2024b). LLM-based metrics are often general enough to evaluate diverse NLG tasks and can provide high evaluation performance without the need for reference texts Liu et al. (2023). They also offer better adaptability to evaluate specific aspects of generated text and are able to evaluate beyond semantic correctness, taking into account aspects like factual consistency or relevance Gu et al. (2025). However, existing LLM-based NLG metrics lack two important features: (1) they are not fully open-source, with most relying on data generated by proprietary LLMs such as GPT-4 OpenAI et al. (2024), and (2) they do not provide precise explanations for their evaluations, with most of them producing only overall scores, or explanations limited to a single short comment (see Table 1 for details).
@@
In this paper, we introduceOpeNLGauge – a versatile, reference-free metric for NLG tasks that provides precise, error-span-based explanations (Figure 1). Unlike existing LLM-based metrics, OpeNLGauge is built entirely on open-weight models and does not rely on annotated datasets. It supports a wide range of NLG tasks, including data-to-text, summarization or story generation. Moreover, it allows for fine-grained evaluation of a specific, customizable aspect of the generated text like faithfulness, fluency or coherence. Our contributions are as follows:
@@
  • We propose an effective prompting strategy and a two-stage LLM ensemble to obtain high-quality evaluations of NLG outputs. Interestingly, the proposed OpeNLGaugeens demonstrates a higher correlation with human judgements on some NLG tasks than previously proposed metrics based on proprietary LLMs.
@@
  • We collect outputs from over a hundred different NLG systems on 15 NLG tasks and use OpeNLGaugeens to annotate them to construct a comprehensive dataset for training open NLG metrics.
@@
  • We use the collected dataset to construct OpeNLGaugeft – a fine-tuned version of the smallest model from the Llama 3.1 family Grattafiori et al. (2024), which is able to provide accurate explanations and good quality assessments of NLG outputs in a more cost-effective way.
@@
  • We perform an extensive meta-evaluation of the proposed metrics on seven datasets covering five NLG tasks. The experimental analysis includes human evaluation of explanation quality, several ablation experiments, and evaluation on tasks, domains and evaluation aspects not seen during training.
@@
Our experiments show that OpeNLGauge achieves higher explanation quality than other metrics trained on data generated by proprietary LLMs, while also delivering strong evaluation performance. On tasks such as summarization, OpeNLGauge achieves higher correlations with human judgments than the strong proprietary GPT-4 model. While the state-of-the-art metric, Themis – which leverages both human-annotated data and data from proprietary LLMs – remains superior overall, OpeNLGauge proves highly competitive and even outperforms it on certain tasks, while also providing more detailed error explanations.
@@
All our experiments were conducted using only quantized open-weight models and two GPUs with 48GB of VRAM, ensuring that the results can be reproduced in many AI research labs. The source code and data are available on Github. [^1] https://github.com/ivankartac/OpeNLGauge
@@ 2 Related Work
Metric No Ref. Aspects Open Err-Span
@@
[…]
Table 1: Comparison of properties in popular LLM-based metrics for NLG: reference-free (No Ref.), customizable aspects (Aspects), evaluation and data generation exclusively using open-weight LLMs (Open), and producing precise explanations with error-span annotation (Err-Span). The compared metrics are: G-Eval Liu et al. (2023), Prometheus (Kim et al., 2024b), Auto-J (Li et al., 2023a), InstructScore (Xu et al., 2023b), TIGERScore (Jiang et al., 2023), Themis (Hu et al., 2024b) and OpeNLGauge (this work).
Context: [SUMMARY OF THE STORY] Question: What is Dr. Heywood Floyd’s mission on the Clavius Base? Answer: According to the summary provided, Dr. Heywood Floyd’s mission on the Clavius Base is to investigate a recently found artifact buried four million years ago. Specifically, his mission is to ride in a Moonbus to the artifact, which is a monolith identical to the one encountered by the man-apes in the summary. Evaluation Aspect: Conciseness – Extent to which the answer is concise and to the point Explanation: This introductory phrase is an error affecting conciseness as it adds unnecessary words to the answer. The response could directly state Floyd’s mission without referencing the source (summary), given the context of the question. Severity: 2 Explanation: This phrase is considered an error affecting conciseness because it provides additional, unnecessary details about the mission. The initial sentence already clearly states Floyd’s mission, making the subsequent elaboration redundant and wordy. Severity: 3 Overall Score: Fair (3/5) Explanation of the Score: While the generated answer generally addresses the question, its conciseness is compromised by the inclusion of redundant information and an unnecessary introductory phrase. Removing these elements would enhance the answer’s directness and efficiency, potentially elevating the score to ‘Good’.
@@
Although NLG has traditionally been evaluated using simple word-overlap-based metrics such as BLEU, these are known to have low correlations with human judgments Novikova et al. (2017a); Reiter (2018). This improved somewhat with the use of trained models for metrics in the past few years Yuan et al. (2021); Zhong et al. (2022); Mehri and Eskenazi (2020b), but the correlations remained moderate. Recently, numerous studies explored the application of LLMs in NLG evaluation. A prominent line of research focuses on leveraging proprietary LLMs such as GPT-4 OpenAI et al. (2024), with direct prompting for an overall score Fu et al. (2024); Kocmi and Federmann (2023b) or even annotating error spans with categories, which has been explored in machine translation Kocmi and Federmann (2023a); Fernandes et al. (2023); Lu et al. (2023). However, using proprietary models is costly and comes with a reproducibility penalty, as some LLM versions become unavailable or are modified in a non-transparent way (Chen et al., 2024). Another concern is data leakage, which affects results, but cannot be controlled in the case of proprietary models Balloccu et al. (2024).
@@
Many recent LLM-based metrics, such as Themis (Hu et al., 2024b), TIGERScore (Jiang et al., 2023), InstructScore Xu et al. (2023b) or Prometheus Kim et al. (2024b), are built on open LLMs, but in fact they all rely on proprietary LLMs such as GPT-4 OpenAI et al. (2024) to generate, filter or annotate training data. Therefore, some of these metrics can be viewed as sophisticated knowledge distillation methods from proprietary to open-weight LLMs. This retains the reproducibility disadvantage, as reconstructing the metric from scratch or adapting it to a new task requires access to closed-source LLMs.
@@
Although several metrics based on open-weight LLMs provide some level of interpretability or explainability, this is often limited to a short, free text review of the output text Hu et al. (2024b); Kim et al. (2024b). However, fine-grained error-span annotation offers several advantages over singular scores and comments. First, it can be easily processed automatically, allowing it to be used in post-processing steps or to provide feedback to the model or training algorithms. Second, it offers greater precision and clarity, as errors are directly associated with a particular part of the output, making it easier to identify and correct issues. Finally, it is more human-like, resembling the output of human annotation schemes such as MQM (Freitag et al., 2021) or ESA (Kocmi et al., 2024). This alignment with human evaluation makes it easier to compare metric-based assessments with human annotations and even allows the metric to serve as an initial annotator, helping to accelerate the human annotation process. While some metrics do provide annotation on the error level, these tend to use closed LLMs and their scope is generally limited to single NLG tasks, such as the machine translation approaches mentioned above or Kasner and Dusek (2024)’s work on data-to-text generation.
@@ 3 Problem Statement
We formulate the problem of evaluating the output of NLG system while providing error-based explanations as follows: Given an input $x$, an output $y$, and an evaluation aspect $a$, the task is to return a tuple $\langle z,\{e_{1},...,e_{n}\}\rangle$, where $z$ is a numeric score assigned to $y$, and $\{e_{1},...,e_{n}\}$ represents a set of error annotations. Each error annotation $e_{i}=\langle s_{i},t_{i},l_{i}\rangle$ includes a span of text $s_{i}\in y$ corresponding to the problematic segment, a textual explanation $t_{i}$, and a severity level $l_{i}$. Although the term error is used throughout this work, it should be understood in a broader sense as any issue in the text related to the evaluated aspect $a$. Examples of outputs provided by OpeNLGauge are presented in Figure 2 and in Appendix H.
@@ 4 Open LLM Ensemble as Evaluator
To achieve a high-quality evaluation of NLG output, we propose a two-stage ensemble of open-weight LLMs. The ensemble consists of $n$ annotator models, which perform independent analyses of the provided NLG output, and a supervisor model, which is responsible for merging their results and filtering inaccuracies. The overview of the approach is presented in Figure 1.
Although our approach requires multiple LLMs, using the ensemble with a handful of LLMs ($n=5$ in our experiments) is still less computationally demanding than running multiple generations to obtain statistical estimates, as required by some metrics, e.g. G-Eval ($n=20$, Liu et al. (2023)). Furthermore, we only use quantized LLMs to limit the computational requirements.
@@
The annotator models are open-weight LLMs that are prompted to identify error spans in the text and to provide detailed explanations as well as severity levels for each error. This approach not only facilitates the interpretability of the evaluation process, but also acts as a chain-of-thought mechanism Wei et al. (2022), helping the model to ground its decisions in a structured reasoning path.
@@
The full annotator model prompt is provided in Appendix E. It contains the description of the evaluated task (e.g., data-to-text), the definition of the evaluated aspect, and a template for the model’s response. Since LLMs are known to confuse different evaluation aspects (Hu et al., 2024a), the prompt also contains several rules that instruct the model to remain focused on the specific evaluation aspect, not to make additional assumptions, and to justify any score lower than the maximum by at least one identified error. Inspired by Liu et al. (2023), we also include detailed steps for error identification.
@@
Furthermore, we provide a description of the overall scoring scale, including an explanation of the lowest and highest scores. The scale is presented as categorical (Unacceptable, Poor, Fair, Good, Excellent), based on the intuition that adjectival categorical scales may be easier for language models to interpret than numerical scales. We use a numerical severity scale (1-5) for scoring individual errors in order to avoid confusion with the overall scoring scale. [^2] A categorical scale was also initially used for error severity. However, we found that LLMs sometimes confused the error severity scale with the overall score scale.
@@
Finally, the prompt contains the input that was used to generate the evaluated output. Some tasks may involve multiple inputs; for instance, evaluating knowledge-grounded question answering requires both question and context knowledge. In such cases, the context is also presented to the model under separate headers. Although structured formats like JSON might make parsing of the output easier, prompting LLMs to reason within strict structured outputs has been shown to impair their performance (Tam et al., 2024; Beurer-Kellner et al., 2024). Therefore, the models are instructed to produce textual outputs.
Supervisor model
@@
The final score of the NLG output is computed by OpeNLGaugeens as a simple average of the scores provided by the annotator models. However, to meet the requirement of explainable output, the error analyses of multiple annotators need to be unified. This is the task of the supervised LMM model.
@@
This open-weight LLM is instructed to: (1) merge errors detected by multiple models, (2) unify their output format, and (3) clean up the annotations. Specifically, the model is instructed to merge all error annotations that refer to the same issue at approximately the same location in the text, while maintaining annotation granularity. It is also prompted to fix potential deviations from the expected output format and to remove any markdown formatting added by the annotators. Finally, the supervisor should clean up outputs of annotator models that occasionally include two non-adjacent spans under a single annotation, or insert comments about the span directly into the location line. To simplify the cleaning process, error analyses produced by outlier annotator models [^3] The annotation is considered to be an outlier if the score provided by the annotator differs from the mean score by at least two standard deviations, and this difference is at least 1. are filtered out from the input to the supervisor model. The full supervisor model prompt is provided in Appendix E.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated great potential as evaluators of NLG systems, allowing for high-quality, reference-free, and multi-aspect assessments. However, existing LLM-based metrics suffer from two major drawbacks: reliance on proprietary models to generate training data or perform evaluations, and a lack of fine-grained, explanatory feedback. We introduce OpeNLGauge, a fully open-source, reference-free NLG evaluation metric that provides accurate explanations based on individual error spans. OpeNLGauge is available as a two-stage ensemble of larger open-weight LLMs, or as a small fine-tuned evaluation model, with confirmed generalizability to unseen tasks, domains and aspects. Our extensive meta-evaluation shows that OpeNLGauge achieves competitive correlation with human judgments, outperforming state-of-the-art models on certain tasks while maintaining full reproducibility and providing explanations more than twice as accurate.
@@
Ivan Kartáč  and  Mateusz Lango  and  Ondřej Dušek Charles University, Faculty of Mathematics and Physics Institute of Formal and Applied Linguistics Prague, Czechia {kartac,lango,odusek}@ufal.mff.cuni.cz
@@ 1 Introduction
Evaluating Natural Language Generation (NLG) systems remains a challenging research problem. Traditional overlap-based metrics, such as BLEU papineni2002bleu and ROUGE lin-2004-rouge, are still widely used but exhibit limited correlation with human judgments, particularly when assessing modern NLG systems novikova-etal-2017-need. With the rise of pre-trained language models, the research community began to shift toward model-based metrics that better capture semantic similarity, yet their performance was still unsatisfactory yan-etal-2023-bleurt; glushkova-etal-2023-bleu.
Figure 1: The ensemble metric OpeNLGaugeens and its distilled version OpeNLGaugeft.
@@
Recently, Large Language Models (LLMs) have demonstrated remarkable potential in imitating human evaluation of generated text jiang2023tigerscore; xu-etal-2023-instructscore; hu2024themis. LLM-based metrics are often general enough to evaluate diverse NLG tasks and can provide high evaluation performance without the need for reference texts liu-etal-2023-g. They also offer better adaptability to evaluate specific aspects of generated text and are able to evaluate beyond semantic correctness, taking into account aspects such as factual consistency or relevance gu2025surveyllmasajudge. However, existing LLM-based NLG metrics lack two important features: (1) they are not fully open-source, with most relying on data generated by proprietary LLMs such as GPT-4 openai2024gpt4technicalreport, and (2) they do not provide precise explanations for their evaluations, with most of them producing only overall scores, or explanations limited to a single short comment (see Table 1 for details).
@@
In this paper, we introduce OpeNLGauge – a versatile, reference-free metric for NLG tasks that provides precise, error-span-based explanations (Figure 1). Unlike existing LLM-based metrics, OpeNLGauge is built entirely on open-weight models and does not rely on human-annotated datasets. It supports a wide range of NLG tasks, including data-to-text, summarization or story generation. Moreover, it allows for fine-grained evaluation of specific, customizable aspects of the generated text such as faithfulness, fluency or coherence. It can also evaluate other user-defined aspects appropriate for a given application.
Our contributions are as follows:
@@
  • We introduce an effective prompting strategy and a two-stage LLM ensemble to obtain high-quality evaluations of NLG outputs. The proposed OpeNLGaugeens demonstrates higher correlations with human judgments on some NLG tasks than previously proposed metrics based on proprietary LLMs.
@@
  • We collect outputs from over 100 NLG systems on 15 NLG tasks and use OpeNLGaugeens to annotate them to construct a comprehensive dataset for training open NLG metrics.
@@
  • The collected dataset is used to construct OpeNLGaugeft – a fine-tuned version of the smallest model from the Llama 3.1 family grattafiori2024llama3herdmodels, which is able to provide accurate explanations and reliable quality assessments of NLG outputs in a more cost-effective way.
@@
  • We perform an extensive meta-evaluation of the proposed metrics on seven datasets covering five NLG tasks. The experimental analysis includes human evaluation of explanation quality, comparison of different score aggregation methods, ablation experiments, and evaluation on tasks, domains, and aspects not seen during training.
@@
Our experiments show that OpeNLGauge achieves higher explanation quality than other metrics trained on data generated by proprietary LLMs, while also delivering strong evaluation performance. On tasks such as summarization, OpeNLGauge achieves higher correlations with human judgments than metrics based on the strong proprietary GPT-4 model. While the state-of-the-art metric, Themis – which leverages both human-annotated data and data from proprietary LLMs – remains superior in terms of correlations with human judgments, OpeNLGauge proves highly competitive and even outperforms it on certain tasks, while also providing more detailed error explanations.
@@
All our experiments were conducted using only quantized open-weight models and two GPUs with 48GB of VRAM each, ensuring that the results can be reproduced in many AI research labs. Our code and data are available on Github. [^1] https://github.com/ivankartac/OpeNLGauge
@@ 2 Related Work
Metric No Ref. Aspects Open Err. Span
@@
[0.5pt/2pt] OpeNLGauge
Table 1: Comparison of properties in different LLM-based metrics for NLG: reference-free (No Ref.), customizable aspects (Aspects), built exclusively using open-weight LLMs (Open), and producing precise explanations with error-span annotation (Err. Span). The metrics compared are: G-Eval liu-etal-2023-g, Prometheus (kim2023prometheus), Auto-J (li2023generative), InstructScore (xu-etal-2023-instructscore), TIGERScore (jiang2023tigerscore), Themis (hu2024themis) and OpeNLGauge (this work).
Context: [SUMMARY OF THE STORY] Question: What is Dr. Heywood Floyd’s mission on the Clavius Base? Answer: According to the summary provided[1], Dr. Heywood Floyd’s mission on the Clavius Base is to investigate a recently found artifact buried four million years ago. Specifically, his mission is to ride in a Moonbus to the artifact, which is a monolith identical to the one encountered by the man-apes in the summary.[2] Evaluation aspect:Conciseness – Extent to which the answer is concise and to the point [1] Explanation: This introductory phrase is an error affecting conciseness as it adds unnecessary words to the answer. The response could directly state Floyd’s mission without referencing the source (summary), given the context of the question. Severity: 2 [2] Explanation: This phrase is considered an error affecting conciseness because it provides additional, unnecessary details about the mission. The initial sentence already clearly states Floyd’s mission, making the subsequent elaboration redundant and wordy. Severity: 3 Overall score: Fair (3/5) Explanation of the score: While the generated answer generally addresses the question, its conciseness is compromised by the inclusion of redundant information and an unnecessary introductory phrase. Removing these elements would enhance the answer’s directness and efficiency, potentially elevating the score to ‘Good’.
@@
Although NLG has traditionally been evaluated using simple word-overlap-based metrics such as BLEU, these are known to have low correlations with human judgments novikova-etal-2017-need; reiter2018structured. This improved somewhat with the use of trained models for metrics in the past few years yuan2021bartscore; zhong2022towards; mehri2020usr, but the correlations remained moderate. Recently, numerous studies explored the application of LLMs in NLG evaluation. A prominent line of research focuses on leveraging proprietary LLMs such as GPT-4 openai2024gpt4technicalreport, with direct prompting for an overall score fu2024gptscore; kocmi2023large or even annotating error spans with categories, which has been explored in machine translation kocmi2023gemba; fernandes2023devil; lu2023error. However, using proprietary models is costly and comes with a reproducibility penalty, as some LLM versions become unavailable or are modified in a non-transparent way (Chen2024How). Another concern is data leakage, which affects results, but cannot be controlled in the case of proprietary models balloccu2024leak.
@@
Many recent LLM-based metrics, such as Themis (hu2024themis), TIGERScore (jiang2023tigerscore), InstructScore xu-etal-2023-instructscore or Prometheus kim2023prometheus, are built on open LLMs, but in fact they all rely on proprietary LLMs such as GPT-4 openai2024gpt4technicalreport to generate, filter, or annotate their training data. Therefore, some of these metrics can be viewed as sophisticated knowledge distillation methods from proprietary to open-weight LLMs. This retains the reproducibility disadvantage, as reconstructing the metric from scratch or adapting it to a new task requires access to closed-source LLMs.
@@
While several metrics based on open-weight LLMs provide some level of interpretability, this is often limited to a short, free-text review of the evaluated output hu2024themis; kim2023prometheus. Fine-grained error-span annotation offers several advantages over singular scores and comments: (1) it can be easily processed automatically, allowing its use in post-processing steps or to provide feedback to the model or training algorithms, (2) it offers greater precision and clarity, as errors are associated with a particular part of the output, making it easier to find and correct issues, (3) it is more human-like, resembling the output of human annotation schemes such as MQM (10.1162/tacl_a_00437) or ESA (kocmi2024error). The alignment with human evaluation allows easier comparison to humans or even use as pre-annotation, helping to accelerate human annotation process. While some metrics do provide annotation on the error level, these tend to use closed LLMs and their scope is generally limited to single NLG tasks, such as the machine translation approaches mentioned above or kasner2024beyond’s work on data-to-text generation.
Scoring outputs with a single LLM may introduce bias (zheng2023judging; panickssery2024llm), which can be alleviated by combining multiple LLMs as evaluators. verga2024replacingjudgesjuriesevaluating apply an ensemble of mostly proprietary LLMs, aggregating their scores either by majority voting over binary ratings, or by averaging for ordinal scores. For pairwise evaluation, li2024prdpeerrankdiscussion propose two methods which combine preferences of multiple LLMs, including an iterative multi-agent discussion.
@@ 3 Problem Statement
We formulate the problem of evaluating the output of an NLG system while providing error-based explanations as follows. Given an input $x$, an output $y$, and an evaluation aspect $a$, the task is to return a tuple $\langle z,\{e_{1},...,e_{n}\}\rangle$, where $z$ is a numeric score assigned to $y$, and $\{e_{1},...,e_{n}\}$ represents a set of error annotations. Each error annotation $e_{i}=\langle s_{i},t_{i},l_{i}\rangle$ includes a span of text $s_{i}\in y$ corresponding to the problematic segment, a textual explanation $t_{i}$, and a severity level $l_{i}$. Although the term error is used throughout this work, it should be understood in a broader sense as any issue in the text related to the evaluated aspect $a$. Examples of outputs provided by OpeNLGauge are presented in Figure 2 and in Appendix L.
@@ 4 Open LLM Ensemble as Evaluator
To achieve a high-quality evaluation of NLG outputs, we propose OpeNLGaugeens, a two-stage ensemble of open-weight LLMs. The ensemble consists of $n$ annotator models, which perform independent analyses of the provided NLG output, and a consolidator model, which is responsible for merging their results and filtering inaccuracies. A high-level overview of the approach is presented in Figure 1.
Although our approach requires multiple LLMs, using the ensemble with a handful of models ($n=5$ in our experiments) is still less computationally demanding than sampling multiple outputs to obtain statistical estimates, as required by some metrics, e.g. G-Eval ($n=20$, liu-etal-2023-g). Furthermore, we only use quantized LLMs to limit the computational requirements.
@@
The annotator models are open-weight LLMs prompted to identify error spans in the text and to provide detailed explanations and severity levels for each error. This approach facilitates the interpretability of the evaluation process, but also acts as a chain-of-thought mechanism 10.5555/3600270.3602070, helping the model to ground its decisions in a structured reasoning path.
@@
The full annotator model prompt is provided in Appendix F. It contains the description of the evaluated task (e.g., data-to-text), the definition of the evaluated aspect, and a template for the model’s response. Since LLMs are known to confuse different evaluation aspects (hu-etal-2024-llm), the prompt also contains several rules that instruct the model to remain focused on the specific evaluation aspect, not to make additional assumptions, and to justify any score lower than the maximum by at least one identified error. Inspired by liu-etal-2023-g, we also include detailed steps for error identification.
@@
Furthermore, we provide a description of the overall scoring scale, including an explanation of the lowest and highest scores. The scale is presented as categorical (Unacceptable < Poor < Fair < Good < Excellent), based on the intuition that adjectival categorical scales may be easier for language models to interpret than numerical scales. We use an integer severity scale (1-5) for scoring individual errors in order to avoid confusion with the overall scoring scale. [^2] A categorical scale was also initially used for error severity. However, we found that LLMs sometimes confused the error severity scale with the overall score scale.
@@
Finally, the prompt contains the input that was originally used to generate the evaluated output (e.g., the source text for summarization or the input data in data-to-text). Some tasks may involve multiple inputs; for instance, evaluating knowledge-grounded question answering requires knowledge of both the question and the context. In such cases, the context is also presented to the model under separate headers. Although structured formats like JSON might make parsing of the output easier, prompting LLMs to reason within strict structured outputs has been shown to impair their performance (tam-etal-2024-speak; kellner2024guiding). Therefore, the models are instructed to produce textual outputs.
Consolidator model
@@
The final score of the NLG output is computed as a simple average of the scores provided by the annotator models. However, to meet the requirement of explainable output, the error analyses of multiple annotators need to be unified. This is the task of the consolidator LLM.
@@
This open-weight LLM is instructed to: (1) merge errors detected by multiple models, (2) unify their output format, and (3) clean up the annotations. Specifically, the model is instructed to merge all error annotations that refer to the same issue at approximately the same location in the text, while maintaining annotation granularity. It is also prompted to fix potential deviations from the expected output format. To simplify the cleaning process, error analyses produced by outlier annotator models [^3] The annotation is considered to be an outlier if the score provided by the annotator differs from the mean score by at least two standard deviations, and this difference is at least 1. are filtered out from the input to the consolidator model. The full prompt for the consolidator model is provided in Appendix F.